VEGAS Team Meeting

Attendees: Ray, Joe, Jason, Melinda, Mark, John, Richard, Paul, Dan, (Steve)

Agenda:

  1. ADC Calibration
  2. Review of “mode 1 ripple” fix / what further tests are required.
  3. Status of “1ms integration time” problem (should have been 0.5 ms).
  4. Status of blanking time / mixed mode discussion and implementation.
  5. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues
  6. Plans for testing for this week.

Meeting notes:

  1. progress on ADC calibration.
    • Jason tested the most recent version of Hong's ADC cal scripts and they work great. If test time can be obtained this week, the next step is to do the ADC cal procedure to the VEGAS roaches. If there's discussion about dividing up test time, and the ADC cal work has a high enough priority, a few hours of time is needed.
    • 9a to 12p opening should be available
  2. Review of “mode 1 ripple” fix / what further tests are required.
    • A new bof file should correct the likely cause. The new file has been installed and then repeat the previous test. Scripts should be available for re-use.
    • A quick test of Mode 2 & 3 should be added if time permits to confirm that there is no common interaction
  3. Status of “1ms integration time” problem (should have been 0.5 ms).
    • Needs to be addressed in Manager. Spec is correct, ("fftrates in S/W wiki). Joe to investigate.
  4. Status of blanking time / mixed mode discussion and implementation.
    • Work has started and unit tests will be used to initially verify the fix - needs a system test to determine how to test on the system (repeat of tsys tests?).
  5. Status of other “current issues”:
  6. galactic center pulsar data has been identified (agbt) - many tvegas directories exist. Possibly simulator data that can be treated as test data - Amanda sjould move forward on file purging, including tvegas.
  7. Plans for testing for this week.
    • Tuesday 8a - 12p - ADC calibrations - Jason
    • Tuesday 9a - 12p - Converter rack - Galen
    • Tuesday 12p - 2p- Paul - Normalize spectra and number of integrations per scan (modes 14 & 19)
    • Tuesday 2p - 3p Anish's set values/blanking time test
    • Tuesday 3p - 4p Richard to test Mode 1 ripple
    • Thursday - TBD Wednesday

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.

- Software engineers should be using a GBT ops software CM&E account

Account 893252

- Electronic engineers should be using a GBT ops electronics CM&E account

Account 432510

- Commissioning activities should be getting charged to hardware support and testing (432510)

Account 432510

- scientists assisting with shared-risk commissioning should be charging an observer support account.

432540 for the most part. Accounts 142023 (GB-based staff) and 142013 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-04-14
Topic revision: r2 - 2014-04-14, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback